Trust in Photographic Presentation

Matthew Gray

April 8, 1997

Introduction

Photography has long been characterized as a medium that does not lie. While framing has always been an issue with photos, it has generally been treated as though a photorealistic image must represent, unchanged, an image of reality.

For the vast majority of photographs, stills, film and video, this is the case. Even more so, this perception is enhanced because the technology typically available to the average person has enforced this. That is, pointand-shoot cameras, video cameras and the like do present a fairly factual representation of reality. For many popular media sources, this is not the case however. As technology improves, it ceases to be the case for the average person as well.

The Problem

As it becomes apparent that images are not always true reflections of reality, it is necassary to have some way of evaluating their veracity. Most people do not have the time, mental energy, or capability to effectively verify the nature of even a small fraction of the images they are regularly presented.

One alternative would be to ignore these images altogether. This might be effective if any substantial fraction of the images encountered were regularly found to be false, but since the majority are actually fair reflections of reality, most are hesitant to discard them wholesale.

Another approach, sometimes consciously used, more often naively used, is to treat all images as real, and hope that altered images either don’t exist, or can be safely ignored. This reasoning, though common, has some serious problems with the increasing ubiquity of such dubious images.

Trust

The solution to this problem is the obvious one. As with other deformable media, such as the spoken word, witten word, paintings and illustrations, the veracity of the content is judged based on trust in the source and the context.

While in many ways, this is obvious, there are some subtleties in the case of images. In general, the ammount of trust assigned to a source is not only a function of the trust in the source and the context, but also affected by the perceived effort required to perform the deceit. That is, an untrusted source would be expected to lie verbally, because it is easy to do so. Lying with an image, however, would be less expected unless there was an easily assignable motive, because creating deceiving images is perceived as difficult (and in the past, has been).

Due to this perception of difficulty in creating a false image, the general behavior is to offer full trust, even if the same source might not be trusted in other, more easily alterable media. Even more so, it is a common beleif that an altered image will be “obvious”, as a blatant lie might be. Further, even obviously retouched or otherwise modified images, are often still considered “true”, if they come from a trusted source, such as certain national media sources.

In general, despite the desire to use a traditional framework of trust to evaluate images, people typically fall into the extreme cases of trusting all images, except the patently obvious deceits (Weekly World News), or distrusting all images (and often all other media).

Conclusion

Photographic (and photorealistic) images have historically been hard to convincingly modify. This history has caused people to approach images with far less skepticism than they approach other media. Digital imaging technologies have so eased the difficulty of modifying images so quickly, that people have not had time to adjust their skepticism. Additionally, these rapid changes in capability have occured faster than people have been able to adjust their patterns of trust. This has resulted in much extremism in trust or lack thereof, in the case of images.